# A Coincidence Algebra Bundle for Decay Quivers: An Algebraic Approach to Gamma-ray Spectroscopy

### v1

@@ 3.1 Coincidence Tensor Space
This larger space contains – as basis vectors – all combinations of paths in the form of tensor products; not only non-connected but overlapping paths as well. However, it is not this total space that we require for our coincidences. In fact, only the physically realized coincidences are needed; these being the pairs of paths that can be observed in a single decay event. For this reason, we consider a subset of this tensor space:
Definition 3.1.
Let $\mathbb{P}(\mathcal{P)}$ be the power set of $\mathcal{P}$, that is, the set of all possible subsets. The coincidence set $\mathcal{O}$ is the subset of $\mathbb{P}(\mathcal{P)}$ and is defined as
$\mathcal{O}\subset\mathbb{P}(\mathcal{P})=\left\{u\in\mathbb{P}(\mathcal{P)}\Big||u|\geq 2,\forall\widehat{p}_{1},\widehat{p}_{2}\in u,s(\widehat{p}_{2})\preceq t(\widehat{p}_{1})\vee s(\widehat{p}_{1})\preceq t(\widehat{p}_{2})\right\},$ (31)
where the paths in a given element of $\mathbb{P}(\mathcal{P)}$ with cardinality greater than or equal to two do not overlap.
@@
Let $o\in\mathcal{O}$ be totally ordered such that $\forall\widehat{p}_{1},\widehat{p}_{2}\in o,\widehat{p}_{1}<\widehat{p}_{2}\iff N(s(\widehat{p}_{2}))<N(s(\widehat{p}_{1}))$; that is, it is ordered by the higher paths in the quiver, and let $o$ be indexed by this ordering. The coincidence vector space is defined as
$\mathcal{C}\subset T(K\mathcal{D}):=span\left\{t=\bigotimes_{i=1}^{|c|}\widehat{p}_{i}\Big|\widehat{p}_{i}\in c,\forall c\in\mathcal{O}\oplus\mathcal{P}\right\}.$ (32)
3.2 Coincidence Algebra Bundle
@@
With the coincidence decay space now defined, we may consider an algebra on this vector space. The path algebra defined for a quiver uses concatenation of paths. In this manner, a given vector, when squared, will acquire components of concatenated paths where the coefficients are multiplied. In the context of decays this alludes to coincidence probabilities or joint probabilities of the two paths. When considering non-connected pairs/sets-of-paths, one must consider what type of multiplication is appropriate when building coincidences out of paths. In the event that the coincidences contain connected paths, it must also reduce to the original path algebra. When considering coincidences and their probabilities, what matters is precisely how they are not connected. A degree of connectivity can be defined by a sum over connecting paths. This is made clear by considering two gammas that are not connected in the decay. If one gamma is at a higher level, what determines if that decay will lead to the second gamma is the sum of the probabilities contained in the different paths that connect them. If the two gammas are directly connected, i.e., concatenate to a larger path, then the sum will simply reduce to the product of their two transition probabilities. However, in the path algebra formalism, the decay scheme exist as a decay vector in the decay vector space; we simply get correct product of probabilities since the transition probabilities are contained in the components of the decay vector, i.e., they are inherently in the vector multiplication by use of the decay vector. When a decay vector existing in the decay vector space – which is now taken to be a subspace of the coincidence decay space – is squared, the product needs to allow for tensor product terms where the coefficients have the appropriate values; i.e., the sum of the paths connecting them. The vector multiplication in the coincidence decay space therefore must depend on the decay vector; i.e., the product of two basis vectors in the coincidence space needs to give a third vector that has a scalar factor which is a function of the decay vector. An algebra therefore can not be defined on the entire vector space and what is needed is an algebra bundle, i.e., a fiber bundle where the base space is the path algebra of the decay quiver and the fibers are a coincidence algebra defined at every point in the base space. Before we define the algebra, let us start with the notion of connectivity between two paths.
Definition 3.3.
@@
$\widehat{p_{i}}\cdot_{d}\widehat{p_{j}}=\begin{cases}c_{d}(\widehat{p_{i}},\widehat{p_{j}})\widehat{p}_{i}\otimes\widehat{p}_{j}&\text{if }s(\widehat{p}_{i})\prec t(\widehat{p}_{j}),\\ \widehat{p_{i}p_{j}}&\text{if }t(\widehat{p}_{i})=s(\widehat{p}_{j}),\\ 0&else.\end{cases}$ (34)
When considering a decays vectors coincidence probabilities that are calculated in its fibre, we require sections that lift that decay vector into its fibre in a manner that copies the vector into the space of the fibre – which is always allowed given $k\mathcal{D}\subset\mathcal{C}$. We then denote the lifting section $l(d):k\mathcal{D}\rightarrow\mathcal{A}_{d}$ such that $d\mapsto d$, and abbreviate to $\tilde{d}:=l(d)$. With the connection between levels and transitions now defined within the algebra, the probability vector of a given transition $\Gamma(d)$ and for two transitions in coincidence $\Gamma_{C}(d)$ can be given by
$\Gamma(d)=\tilde{b}\cdot_{d}\tilde{\tau}$, (35)
$\Gamma_{C}=\tilde{b}\cdot_{d}\tilde{\tau}\cdot_{d}\tilde{\tau}$. (36)
@@
As with the path algebra, the order of multiplication is important and must be ordered in the hierarchy of the quiver, that is, the first term contains the branching ratios on the stationary paths which is connected to all paths below it via the scalar connection defined in the coincidence algebra. To get the coincidence probabilities when then further connect it to all paths below it – this is extract by simply coincidence multiplying onto the (lifted) transition vector again. What is different now from the path algebra formulation, as we will see shortly, is that we now have terms in the probability vector that correspond to coincidences of paths that are not directly connected; e.g., with this a term for the ”summing in” of $x_{32}$ and $x_{10}$ is now given. Not only that, but for each transition, the level at which it decay from is now a separate basis vector as the transition is now tensored with the vertices. Let us now consider how detection maps work in the base space on the coincidence algebra bundle.
3.3 Detection maps in the bundle.
@@
With the detection maps taking a decay vector to a new vector in the path algebra that alters the probabilities to account for detection effects, when we calculate coincidence probabilities in that detection mapped fibre, the probabilities now alter the coincidence multiplication. When dealing with transporting vectors across a bundle one must also consider the curvature that may be present in the bundle and whether a vector can trivially be mapped as usual. A connection444separate from the scalar connection defined above, though the scalar connection would go into defining the multiplicative connection when analyzing the curvature of the algebra across the bundle is usual defined on a bundle that allows one to calculate how sections can be moved to different fibres. For this present work, we do not yet concern ourselves with explicitly calculating the connection on a coincidence algebra bundle. However, we do note that the underlying vector bundle on which the fibrewise algebra is defined is a product bundle and thus admits a trivial connection with zero curvature. When considering the multiplicative connection, that is, the connection that is compatible with the coincidence algebra defined on the fibres, one then must be careful and acknowledge that transporting products of vectors is different than taking products of transported vectors – we save this investigation for later work. It is sufficient for now to realize that we may transport a vector in one fibre to another fibre with out having to consider a connection given that a trivial connection can be established on the underlying product vector bundle. This is important when dealing with multiple detection maps to calculate a given probability vector. Whenever we are multiplying paths together, it must be defined within the fibre of a given point in the base space, i.e, it must be defined for a given decay vector. When taking the coincidence product in a detection mapped fibre, the probabilities that connect the two paths in the product will be the ones defined by the detection map. For transitions of interest, a detection map that accounts for full energy peak efficiencies would be desired. However, the probabilities that connect the transitions of interest would be from a detection mapped decay vector that accounts for summing out probabilities. Therefore, we wish to calculate summing out probabilities in the fibre at the summing out mapped decay vector, but use vectors transported from other vectors in the base space. We may therefore have maps between points in the base space, or maps between fibres. Since these maps act only on decay vectors which live in the path algebra which is also a subspace of the coincidence algebra we may either consider lifting the decay maps in the base space or lifting the decay vector and mapping it to the fibre in which probabilities are calculated in. Again, given the trivial product vector bundle underlying the algebra bundle, the manner in which we move vectors around does not matter, it is only the transportation of products that changes. Put more aptly, the detection maps are not homomorphisms of the coincidence algebra defined on the fibres. When considering the power expansion of detection mapped vectors calculated in the coincidence algebra, we must now distinguish what map is being applied to the vector being multiplied and what detection map is defining the coincidence multiplication. when considering the same problem, of coincidence summing with summing out and summing in, we may use the follow power expansion where the fibre is now explicitly stated:
$\overline{\tilde{\varphi}_{h}(\tau)}^{n_{\varphi_{o}(\tau)}}=\sum_{k=1}^{n}\tilde{\varphi}_{h}(\tau)^{k_{\varphi_{o}(\tau)}}=\tilde{\varphi}_{h}(\tau)+\tilde{\varphi}_{h}(\tau)\cdot_{\varphi_{o}(\tau)}\tilde{\varphi}_{h}(\tau)+\tilde{\varphi}_{h}(\tau)\cdot_{\varphi_{o}(\tau)}\tilde{\varphi}_{h}(\tau)\cdot_{\varphi_{o}(\tau)}\tilde{\varphi}_{h}(\tau)+$... (37)

### v2

@@ III.1 Coincidence Tensor Space
This larger space contains – as basis vectors – all combinations of paths in the form of tensor products; not only non-connected but overlapping paths as well. However, it is not this total space that we require for our coincidences. In fact, only the physically realized coincidences are needed–these being the pairs of paths that can be observed in a single decay event. For this reason, we consider a subset of this tensor space:
Definition III.1.
Let $\mathbb{P}(\mathcal{P}_{0}\cup\mathcal{P}_{1})$ be the power set of $\mathcal{P}_{0}\cup\mathcal{P}_{1}$, that is, the set of all possible subsets of $\mathcal{P}_{0}\cup\mathcal{P}_{1}$. The coincidence set $\mathcal{O}$ is a subset of $\mathbb{P}(\mathcal{P}_{0}\cup\mathcal{P}_{1})$ and is defined as
$\mathcal{O}\subset\mathbb{P}(\mathcal{P}_{0}\cup\mathcal{P}_{1})=\left\{u\in\mathbb{P}(\mathcal{P}_{0}\cup\mathcal{P}_{1})\Big||u|\geq 2,\forall\widehat{p}_{1},\widehat{p}_{2}\in u,s(\widehat{p}_{2})\preceq t(\widehat{p}_{1})\vee s(\widehat{p}_{1})\preceq t(\widehat{p}_{2})\right\},$ (29)
where the paths in a given element of $\mathbb{P}(\mathcal{P}_{0}\cup\mathcal{P}_{1})$ with cardinality greater than or equal to two do not overlap.
@@
Let $o\in\mathcal{O}$ be totally ordered such that $\forall\widehat{p}_{1},\widehat{p}_{2}\in o,\widehat{p}_{1}<\widehat{p}_{2}\iff N(s(\widehat{p}_{2}))<N(s(\widehat{p}_{1}))$; that is, it is ordered by the higher paths in the quiver, and let $o$ be indexed by this ordering. The coincidence vector space $\mathcal{C}\subset T(k\mathcal{D})$ is defined as
$\mathcal{C}:=span\left\{t=\bigotimes_{i=1}^{|c|}\widehat{p}_{i}\Big|\widehat{p}_{i}\in c,\forall c\in\mathcal{O}\cup\mathcal{P}_{0}\cup\mathcal{P}_{1}\right\}.$ (30)
III.2 Coincidence Algebra Bundle
@@
With the coincidence decay space now defined, we may consider an algebra on this vector space. The path algebra defined for a quiver uses concatenation of paths. In this manner, a given vector, when squared, will acquire components of concatenated paths where the coefficients are multiplied (i.e., it is a graded algebra). In the context of decays this alludes to coincidence probabilities or joint probabilities of the two paths. When considering non-connected pairs/sets-of-paths, one must consider what type of multiplication is appropriate when building coincidences out of paths. In the event that the coincidences contain connected paths, it must also reduce to the original path algebra. When considering coincidences and their probabilities, what matters is precisely how they are not connected. A degree of connectivity can be defined by a sum over connecting paths. This is made clear by considering two gammas that are not connected in the decay. If one gamma is at a higher level, what determines if that decay will lead to the second gamma is the sum of the probabilities contained in the different paths that connect them. This can be seen in Figure 2 where gamma A and C are only partially connected given that there a exist a decay from A to the ground state that does not emit gamma C. The connectivity between A and C is the probability that the tail vertex of A leads to the source vertex of C given the intermediary transitions.
If the two gammas are directly connected, i.e., concatenate to a larger path, then the sum will simply reduce to the product of their two transition probabilities. However, in the path algebra formalism, the decay scheme exist as a decay vector in the decay vector space; we simply get correct product of probabilities since the transition probabilities are contained in the components of the decay vector, i.e., they are inherently in the vector multiplication by use of the decay vector. When a decay vector existing in the decay vector space – which is now taken to be a subspace of the coincidence decay space – is squared, the product needs to allow for tensor product terms where the coefficients have the appropriate values; i.e., the sum of the paths connecting them. The vector multiplication in the coincidence decay space therefore must depend on the decay vector; i.e., the product of two basis vectors in the coincidence space needs to give a third vector that has a scalar factor which is a function of the decay vector. An algebra therefore can not be defined on the entire vector space and what is needed is an algebra bundle, i.e., a fiber bundle where the base space is the path algebra of the decay quiver and the fibers are a coincidence algebra defined at every point in the base space. Before we define the algebra, let us start with the notion of connectivity between two paths.
Definition III.3.
@@
$\widehat{p_{i}}\cdot_{d}\widehat{p_{j}}=c_{d}(\widehat{p_{i}},\widehat{p_{j}})\widehat{p}_{i}\otimes\widehat{p}_{j}$ (32)
If $\widehat{p}_{i}$ and $\widehat{p}_{j}$ are not directly connected (they do not form a path) but are ordered via Def. III.2, than $0\leq c_{d}(\widehat{p_{i}},\widehat{p_{j}})\leq 1$. If they do form a path than $c_{d}(\widehat{p_{i}},\widehat{p_{j}})=1$. If they overlap than the product returns zero. Note, since we are in a subspace of $\mathcal{T}(k\mathcal{D})$, we notate concatenated arrows, i.e., paths of length greater than one, using tensor product notation instead of the hat notation of the path algebra. We allow this as the coincidence space does not distinguish between the ability to concatenate directly–it only requires coincidence. The coincidence algebra extends the path algebra into a tensor space of coincidences in general, not just paths. When considering a decays vectors coincidence probabilities that are calculated in its fibre, we require sections that lift that decay vector into its fibre in a manner that copies the vector into the space of the fibre – which is always allowed given $k\mathcal{D}\subset\mathcal{C}$. We then denote the lifting section $l(d):k\mathcal{D}\rightarrow\mathcal{A}$ such that $d\mapsto(d,d)$, and abbreviate to $\tilde{d}:=l(d)$. The underlying vector bundle of the algebra bundle is a product bundle $k\mathcal{D}\times\mathcal{C}$, thus sections (vectors in the bundle) are maps into the tuple $(a,b)$ with $a\in k\mathcal{D}$ and $b\in\mathcal{C}$. However, when considering a vector in the bundle, the base space vector will be which defines the algebra; what is being multiplied as a vector is the vector in $\mathcal{C}$,i.e., multiplying two vectors in a fibre defined at point $a\in k\mathcal{D}$ is defined as the multiplication of two sections at that fibre.
$b\cdot_{a}c=(a,b)(a,c)=(a,b\cdot c)$ (33)
With this is mind, further notation restricts to writing the coincidence vector only with the path vector defined in the product. With the connection between levels and transitions now defined within the algebra, the probability vector of a given transition $\Gamma_{1}(d)$ and for two transitions in coincidence $\Gamma_{2}(d)$ can be given by
$\Gamma_{1}(d)=\tilde{b}\cdot_{d}\tilde{\tau}$, (34)
$\Gamma_{2}(d)=\tilde{b}\cdot_{d}\tilde{\tau}\cdot_{d}\tilde{\tau}$. (35)
@@
As with the path algebra, the order of multiplication is important and must be ordered in the hierarchy of the quiver, that is, the first term contains the branching ratios on the stationary paths which is connected to all paths below it via the scalar connection defined in the coincidence algebra. To get the coincidence probabilities we then further connect it to all paths below it; this is done by simply coincidence multiplying onto the (lifted) transition vector again. What is different now from the path algebra formulation is that we now have terms in the probability vector that correspond to coincidences of paths that are not directly connected; e.g., with this a term for the coincidence of $x_{32}$ and $x_{10}$ is now given. Not only that, but for each transition, the level at which it decayed from is now a separate basis vector as the transition is now tensored with the vertices. Examples for $\Gamma_{1}(d)$ and $\Gamma_{2}(d)$ for a general decay scheme with $n=3$ is shown in Tables 1 and 2, respectfully.
Basis Vector Branch $f_{3}$ $f_{2} Branch$ $f_{1}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{30}$ $f_{3}x_{30}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{31}$ $f_{3}x_{31}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{32}$ $f_{3}x_{32}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{21}$ $f_{3}x_{32}x_{21}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{20}$ $f_{3}x_{32}x_{20}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{10}$ $f_{3}(x_{31}+x_{32}x_{21})x_{10}$
$\widehat{\epsilon}_{2}\otimes\widehat{p}_{20}$ $f_{2}x_{20}$
$\widehat{\epsilon}_{2}\otimes\widehat{p}_{10}$ $f_{2}x_{21}x_{10}$
$\widehat{\epsilon}_{1}\otimes\widehat{p}_{10}$ $f_{1}x_{10}$
Table 1: Probability vector $\Gamma_{1}(d)$ coefficients for a 4 level decay quiver. The basis vectors determine the transition and the coefficients give the probability of that transition for a given decay. The basis vectors are separated by the branch they decayed from.
[…]
$\widehat{\epsilon}_{3}\otimes\widehat{p_{31}}\otimes\widehat{p_{10}}$ $f_{3}x_{31}x_{10}$
$\widehat{\epsilon}_{3}\otimes\widehat{p_{32}}\otimes\widehat{p_{20}}$ $f_{3}x_{32}x_{20}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{32}\otimes\widehat{p}_{10}$ $f_{3}x_{32}x_{21}x_{10}$
$\widehat{\epsilon}_{3}\otimes\widehat{p_{32}}\otimes\widehat{p_{21}}$ $f_{3}x_{32}x_{21}$
$\widehat{\epsilon}_{3}\otimes\widehat{p_{21}}\otimes\widehat{p_{10}}$ $f_{3}x_{32}x_{21}x_{10}$
$\widehat{\epsilon}_{2}\otimes\widehat{p_{21}}\otimes\widehat{p_{10}}$ $f_{2}x_{21}x_{10}$
Table 2: Probability vector $\Gamma_{2}(d)$ coefficients for a 4 level decay quiver. The basis vectors determine the coincidence and the coefficients give the probability of that coincidence for a given decay. The basis vectors are separated by the branch they decayed from.
We also point out how easily gated probability vectors emerge out of the coincidence algebra–indeed, that is its strength. The transition matrix formalism allows for the summing corrections for all of the singles spectra (single gamma ray peaks) in the form of a matrix. As shown in equation 6, the coincidence probability for two specific gamma can be given; however, there is no equation the produces all possible coincidences (such would require a rank for tensor, with two indices for each gamma). In [7], they showed how to calculate the final probability matrix with summing corrections for a given gate (gamma in coincidence) by partitioning the transition matrix into block matrices and treating the block matrices with the Semkow formalism separately before direct summing them back together. In the coincidence algebra, all possible coincidences can be given in by single vector. To gate on a certain gamma, $\Gamma_{C}(d)$ can be projected onto a subspace defined by the gate. The partitioned Semkow formalism in [7] indeed gives the transition matrix representation of $\Gamma_{C}(d)$ projected onto a gated subspace. Projections of coincidence vectors can be defined in the following manner. First, we begin by formulating how to compare tensor products of paths. This is can be done by defining forms that sum over the paths in a tensor product:
$\widehat{p}_{i},\widehat{p}_{1}\otimes\widehat{p}_{2}\otimes...\otimes\widehat{p}_{j}]_{p}=\sum_{k=1}^{j}\langle\widehat{p}_{k},\widehat{p_{i}}\rangle_{p},$ (36)
which expands linearly for a given vector $c\in\mathcal{C}$ by
$\widehat{p}_{i},c]_{p}=\sum_{q=1}^{m}a_{q}[\widehat{p}_{i},\widehat{c}_{q}]_{p},$ (37)
where
$c=\sum_{q=1}^{m}a_{q}\widehat{c}_{q}$ (38)
with $\widehat{c}_{q}$ denoting some tensor product of paths (a basis vector in $\mathcal{C}$). We therefore define the following gate projector
Definition III.5.
The gate projector $G(\widehat{p}_{i},\widehat{c}):\mathcal{C}\rightarrow\mathcal{C}$ projects a vector $c\in\mathcal{C}$ onto the subspace whose basis vectors overlap with a path $\widehat{p}_{i}$ such that $\widehat{p}_{i},\widehat{c}]_{p}>0$.
$G_{1}(\widehat{p}_{i},\widehat{c})=\sum_{q=0}^{m}a_{q}[\widehat{p}_{i},\widehat{c}_{q}]_{P}$ (39)
For example, given the coincidence vector in Table 2,
$G(\widehat{p}_{32},\Gamma_{2}(d))=$ $f_{3}x_{32}x_{20}\widehat{p_{32}}\otimes\widehat{p_{20}}+f_{3}x_{32}x_{21}x_{10}\widehat{p}_{32}\otimes\widehat{p}_{10}$ (40)
$+f_{3}x_{32}x_{21}\widehat{p_{32}}\otimes\widehat{p_{21}}$
The Gate projector may be extended to gate on coincidences, not just single transitions:
$G_{2}(\widehat{p}_{i},\widehat{p}_{l},\widehat{c})=\sum_{q=0}^{m}a_{q}[\widehat{p}_{i},\widehat{c}_{q}]_{p}[\widehat{p}_{l},\widehat{c}_{q}]_{p}$ (41)
This can of course easily be extended to higher order coincidences by multiply on a form for each given transition. Let us now consider how detection maps work in the base space on the coincidence algebra bundle.
III.3 Detection maps in the bundle.
@@
With the detection maps taking a decay vector to a new vector in the path algebra that alters the probabilities to account for detection effects, when we calculate coincidence probabilities in that detection mapped fibre, the probabilities now alter the coincidence multiplication. When dealing with transporting vectors across a bundle one must also consider the curvature that may be present in the bundle and whether a vector can trivially be mapped as usual. A connection [^2] This is separate from the scalar connection defined above, though the scalar connection would go into defining the multiplicative connection when defining parallel transport across the bundle is usual defined on a bundle that allows one to calculate how sections can be moved to different fibres. For this present work, we do not yet concern ourselves with explicitly calculating the connection on a coincidence algebra bundle. However, we do note that the underlying vector bundle on which the fibrewise algebra is defined is a product bundle and thus admits a trivial connection with zero curvature. When considering the multiplicative connection, that is, the connection that is compatible with the coincidence algebra defined on the fibres, one then must be careful and acknowledge that transporting products of vectors is different than taking products of transported vectors – we save this investigation for later work. It is sufficient for now to realize that we may transport a vector in one fibre to another fibre with out having to consider a connection given that a trivial connection can be established on the underlying product vector bundle. This is important when dealing with multiple detection maps to calculate a given probability vector. Whenever we are multiplying paths together, it must be defined within the fibre of a given point in the base space, i.e, it must be defined for a given decay vector. When taking the coincidence product in a detection mapped fibre, the probabilities that connect the two paths in the product will be the ones defined by the detection map. For transitions of interest, a detection map that accounts for full energy peak efficiencies would be desired. However, the probabilities that connect the transitions of interest would be from a detection mapped decay vector that accounts for summing out probabilities. Therefore, we wish to calculate summing out probabilities in the fibre at the summing out mapped decay vector, but use vectors transported from other vectors in the base space (the full energy ones). We may therefore have maps between points in the base space, or maps between fibres. Since these maps act only on decay vectors which live in the path algebra which is also a subspace of the coincidence algebra we may either consider lifting the decay maps in the base space or lifting the decay vector and mapping it to the fibre in which probabilities are calculated in. Again, given the trivial product vector bundle underlying the algebra bundle, the manner in which we move vectors around does not matter, it is only the transportation of products that changes. Put more aptly, the detection maps are not homomorphisms of the coincidence algebra defined on the fibres. When considering the power expansion of detection mapped vectors calculated in the coincidence algebra, we must now distinguish what map is being applied to the vector being multiplied and what detection map is defining the coincidence multiplication. when considering the same problem, of coincidence summing with summing out and summing in, we may use the following power expansion where the fibre is now explicitly stated:
$\overline{\tilde{\varphi}_{h}(\tau)}_{N}^{n_{\varphi_{o}(\tau)}}=\sum_{k=1}^{n}\tilde{\varphi}_{h}(\tau)^{k_{\varphi_{o}(\tau)}}=\tilde{\varphi}_{h}(\tau)+\frac{\tilde{\varphi}_{h}(\tau)\cdot_{\varphi_{o}(\tau)}\tilde{\varphi}_{h}(\tau)}{N}+\frac{\tilde{\varphi}_{h}(\tau)\cdot_{\varphi_{o}(\tau)}\tilde{\varphi}_{h}(\tau)\cdot_{\varphi_{o}(\tau)}\tilde{\varphi}_{h}(\tau)}{N^{2}}+$... (42)